# A Regret Perspective on Online Selective Generation

### v1

@@ 4.3 Regret Bound
Let $\ell_{t}(\cdot)\in[0,\ell_{\text{max}}]$. For any $T\in\mathbb{N}$ and finite hypotheses $\mathcal{{H}}$, Algorithm 1 provides the following regret bound if $\eta=\sqrt{{\ln|\mathcal{H}|}\mathbin{/}{\ell_{\text{max}}^{2}T}}$:
$\textbf{Reg}_{T}=\sum_{t=1}^{T}\mathbbm{E}_{\tau,\tau_{t}\sim p_{t}}\ell_{t}(\tau,\alpha\mid\mathcal{{H}}_{t}({\tau_{t}}))-\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\ell_{t}(\tau,\alpha)\leq 2\ell_{\text{max}}\sqrt{T\ln|\mathcal{H}|}$
for any observed loss sequences $\ell_{t}$ generated before learning by an oblivious adversary.
Interestingly, despite of partial feedback, our algorithm achieves the same upper bound as EW with full feedback, i.e., $\ell_{\text{max}}\sqrt{{T\ln|\mathcal{{H}}|}/{2}}$, up to a constant factor due to our unbiased loss estimator (8) with rich information, whereas Exp3 suffers from an additional factor of $\sqrt{|\mathcal{{H}}|}$ (see (3)) due to its unbiased loss estimator with limited information.
4.4 Regret-to-FDR Conversion
We mainly leverage algorithms that minimize the regret $\textbf{Reg}_{T}$ for learning selective generators. Yet, the key requirement for selective generation is the FDR guarantee of a learner at a desired level. To this end, we introduce a novel perspective on the connection of a regret bound to an FDR bound. In particular, any learner, including learners for online learning and bandits, that minimizes the regret (1) controls the FDR (4). This can be achievable as the designed loss (6) is penalized with the regularized parameter $\lambda$ when the FDR loss at each step is higher than $\alpha$ by $d_{t}(\tau,\alpha)$, resulting in controlling the FDR over time $T$ by $\alpha$. The following is a proposed regret-to-FDR conversion lemma. See Appendix G.3 for a proof.
Lemma 1.
Let $T\in\mathbb{N}$, $\alpha\in(0,1)$. If we have a learner that achieves a bounded regret for any loss sequences of (6), then the learner also bounds the FDR, i.e., if $\lambda\geq 1/T^{1/4}$
$\textbf{FDR}_{T}\leq\alpha+\frac{1+\textbf{Reg}_{T}/T}{(1-\textbf{Ineff}_{T})T^{1/4}}.$
Importantly, this lemma is applicable to learning for both full and partial feedback, as it is agnostic to feedback mechanisms. Also, the lemma implies that if $\textbf{Reg}_{T}$ has a sublinear bound, satisfied by most online learners, and $\textbf{Ineff}_{T}$ is constant less than 1, which also mostly holds, then $\textbf{FDR}_{T}$ gets closer to $\alpha$ at most the rate of $\mathcal{{O}}(1/T^{1/4})$, i.e., $\textbf{FDR}_{T}\leq\alpha+\mathcal{{O}}(1/T^{1/4})$. Note that $\textbf{Ineff}_{T}$ does have an effect on the convergence speed by a constant factor. See Section H for discussion on efficiency.
@@ 5 Experiments
We empirically justify that ExSUL controls the FDR while maximizing selection efficiency. Moreover, we demonstrate the controllability guarantee holds under diverse environments. In particular, we consider three environments: (1) stochastic, (2) distribution-shift, and (3) interactive environments.
Datasets and Models. We use two datasets for stochastic and distribution-shift environments, $79\mathrm{K}$ Natural Question (NQ) kwiatkowski2019natural and $93\mathrm{K}$ TriviaQA joshi2017triviaqa with two base models, GPT-3.5-turbo openai2023gpt35turbo and LLaMA3.1-8B-Instruct meta2024llama31-8binstruct . We simulate distribution-shift environments by mixing two datasets in diverse ways and an interactive environment by generating interaction between two GPT-3.5-turbo models (see Section E.2 and E.3, respectively for details). To obtain entailment labels and self-consistency scores, we additionally use GPT-3.5-turbo as an entailment model.
Scoring functions. We consider two scoring functions $f_{\texttt{std}}$ and $f_{\texttt{con}}$, introduced in Section 2.1. In short, $f_{\texttt{std}}$ is a standard log-likelihood score defined as the conditional probability of an answer given a question. $f_{\texttt{con}}$ is a self-consistency score of an answer computed via entailment scores (from an entailment model) across samples. We use $f_{\texttt{con}}$ unless specified.
Methods. We consider two baselines, Exp3-SG and No-SG, and one performance upper bound algorithm, EW-SG. In particular, (1) Exp3-SG (Algorithm 5) is an online selective generator with partial feedback by converting Exp3 auer2002nonstochastic for selective generation via our regret-to-FDR conversion lemma. (2) No-SG is a non-selective generator (i.e., $\tau_{t}=0$ always), to serve as a standard use of a generator without abstention. (3) EW-SG (Algorithm 4) is an online selective generator with full feedback by converting this for selective generation via our conversion lemma. As it exploits full feedback, we use it as our empirical performance upper bound.
Metrics. We measure performance of ours along with comparing methods via the FDR, i.e., $\textbf{FDR}_{t}$ and selection inefficiency, i.e., $\textbf{Ineff}_{t}$. Note that No-SG achieves zero selection inefficiency by definition, so we do not explicitly add in figures.
Parameters. A desired FDR level $\alpha$ and time horizon $T$ depend on tasks, but we set $|\mathcal{{H}}|=1\mathrm{K}$ and $\lambda=T^{\nicefrac{{1}}{{4}}}$ by default. Note that $\alpha$ is a desired FDR set by users. This usually depends on tasks, e.g., $\alpha$ is smaller than an error of a base generator on a given task without selection to avoid a trivial FDR.
@@ 5.1 Stochastic Environment
Figure 2: Comparison of selective generation methods under a stochastic environment with LLaMA3.1-8B-Instruct as a generator on TriviaQA ($T=30\mathrm{K},\alpha=0.08$). The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples over $30$ random trials.
@@
Figure 3: Comparison of selective generation methods under a stochastic environment with GPT-3.5-turbo as a generator on TriviaQA ($T=30\mathrm{K},{\alpha=0.08}$). The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples over $30$ random trials.
In Figure 2 and 3, we observe that ExSUL converges fast under $\alpha$, comparable to EW-SG which has access to full feedback, while Exp3-SG fails to converge under $\alpha$ in most cases. This is because Exp3-SG only observes partial feedback from the chosen arm, requiring a much longer time horizon $T$ for convergence (Figure 23), which supports the regret bound analysis discussed in Section 4.3. These results imply that our partial feedback unlocking is beneficial in a regret bound and faster convergence to the desired FDR. Similar trends are observed in Figure 9 and 11. Additionally, Figure 21 demonstrates that our algorithm effectively controls the FDR under varying $\alpha$ in practice. Note that algorithms with $f_{\texttt{con}}$ have faster convergence than those with $f_{\texttt{std}}$ (e.g., Figure 9 v.s. 11), suggesting that a well-calibrated scoring function can have a positive effect on convergence speed.
5.2 Distribution-shift Environment
In this study, to evaluate the robustness of selective generation we examine three distinct distribution-shift scenarios. First, we create a single instantaneous change by concatenating one large segment from one dataset with an other segment from the other. Second, we induce frequent, periodic shift by interleaving a series of small, fixed-size chunks drawn randomly from each dataset. Third, we provide progressive shift by sampling each example according to a mixing probability that varies linearly. More details of environment setup introduce in Appendix E.2.
In Figure 4,5, and 6, we observes that ExSUL is clearly better in convergence speed than Exp3-SG, while closely follows EW-SG for data distribution shifted environment. Interestingly, in Figure 4, Exp3-SG exhibits a sharp increase in the FDR immediately following the distribution shift, highlighting its sensitivity to sudden changes in data distribution, while the proposed ExSUL does not. In 22, as in the stochastic environments, we also confirmed that our methods find selective generators for a diverse desired FDR $\alpha$, showing theoretical controllability empirically. See additional results on varying generators and shift types in Section F.2 and a longer time horizon in Section 24.
@@ 5.2 Distribution-shift Environment
Figure 4: Comparison of selective generation methods under a distribution-shift environment with GPT-3.5-turbo as a generator ($T=120\mathrm{K},\alpha=0.1$). Here, we consider a single distribution shift from $60\mathrm{K}$-sized TriviaQA to $60\mathrm{K}$-sized NQ, denoted in a dotted vertical line. The violin plots are drawn with randomly chosen $120\mathrm{K}$ samples with $30$ random trials.
@@
Figure 5: Comparison of selective generation methods under a distribution-shift environment with GPT-3.5-turbo as a generator ($T=120\mathrm{K},\alpha=0.1$). We consider multiple distribution shifts by alternating $10\mathrm{K}$-sized TriviaQA and NQ 6 times (i.e., TriviaQA, NQ, TriviaQA, NQ and so forth), where shifting points are visualized in dotted vertical lines. The violin plots are drawn with randomly chosen $120\mathrm{K}$ samples with $30$ random trials.
@@
(c) FDR distribution
Figure 6: Comparison of selective generation methods under a distribution-shift environment with GPT-3.5-turbo as a generator ($T=60\mathrm{K},\alpha=0.1$). We consider a distribution shift by gradually shifting from TriviaQA to NQ over time. The violin plots are drawn with randomly chosen $60\mathrm{K}$ samples with $30$ random trials.
5.3 Interactive Environment
We empirically demonstrate the efficacy of our proposed ExSUL in real-world interactive applications. In particular, we simulate an interactive environment by implementing a user-acting agent, a question-answering agent, and an evaluating agent by states. These agents interact over multiple turns, forming a dialog-like sequence of states that mimics realistic user interactions, and our selective generation method ExSUL decides whether to abstain from generation as shown in Figure 1. In Figure 7, we observe that ExSUL consistently controls the FDR under $\alpha$, despite the instability of $\textbf{FDR}_{t}$ due to the shifting dynamics of question-answering. This observation demonstrates that our method ExSUL is robust under frequently shifting distributions over time. In Figure 1, ExSUL successfully abstains from incorrect answers, helping to achieve the desired FDR level $\alpha$.
(a) FDR over time
(b) Inefficiency over time
Figure 7: $\textbf{FDR}_{t}$ and $\textbf{Ineff}_{t}$ for ExSUL under an interactive environment with GPT-3.5-Turbo as a question-answering agent ($T=15{,}000$, $\alpha=0.25$). A user agent continuously creates questions based on changing context, which simulate distribution shifts.
@@ 6 Conclusion
We propose a learning algorithm for online selective generation with partial feedback, which provides an FDR controllability guarantee, while maximizing selection efficiency. To this end, we leverage the well-developed adversarial bandit problem. In particular, we reduce selective generation to adversarial bandits and extend the Exp3 algorithm for adversarial bandits by leveraging a technique called partial feedback unlocking, which addresses the lack of feedback information due to the nature of partial feedback. Finally, we introduce a novel Regret-to-FDR conversion lemma such that we can use any regret minimization algorithm, including our extended Exp3 for selective generation to control an FDR at a desired level. We theoretically and empirically justify our method by providing an efficient regret bound and evaluating diverse learning environments, including stochastic, distribution-shift, and interactive environments.
Limitations. As for the main limitation of our method, we assume that an oblivious adversary generates data sequences, i.e., a data sequence is drawn before learning, so our FDR guarantee may not hold if an stronger adversary generates data sequences adaptively.
@@ Acknowledgment
Appendix A Additional Related Work
Constrained Online Learning Several works have studied online learning with constraints. [19] first proposes long-term constraints in online convex optimization with time-invariant constraints, which allows to have the constraint function $g_{t}$ to be violated at each time step, i.e., $g(\mathbf{x}_{t})>0$, but eventually satisfied, i.e., $\sum_{t=1}^{T}g(\mathbf{x}_{t})\leq 0$. Some other works solve a similar problem but with time-varying constraints, which is more challenging. [20, 21] solve a network resource allocation task by viewing this as an constrained optimization problem. Especially, [21] solves the problem without a convex assumption on loss. [22] gives analysis on optimal algorithms for online convex optimization with adversarial constraints. We similarly consider a FDR guarantee to be satisfied over time but with partial feedback.
Conformal Prediction with Bandit Feedback There is little work about selective or conformal prediction in bandit settings. [23] proposes an online conformal prediction method but with semi-bandit feedback, where we can observe the true label only if it is predicted. Our method follows the practical motivation in having partial feedback idea but we consider selective generation.
Appendix B Additional Preliminary
B.1 Online Learning and EW Algorithm
Online learning considers to design a learner that adapts to any sequence of full feedback on hypotheses without any or with little assumption on data generation process. In particular, the goal of an online learner is to find a distribution $p_{t}$ over hypotheses with full feedback that minimizes the regret $\textbf{Reg}_{T}$.
Along with many online learning algorithms, we introduce Exponential Weighting (EW) [14], which minimizes the regret with finite hypotheses (Algorithm 2). In particular, EW maintains the goodness of weights for each hypothesis in terms of cumulative loss and updates them from feedback. The feedback of a hypothesis $\tau_{t}$ at time $t$ is represented by loss $\ell_{t}(\tau_{t})$ but we can compute this loss for all hypotheses other than a chosen hypothesis $\tau_{t}$ as we directly observe a true label $\mathbf{y}_{t}$ at the same time $t$. This EW algorithm satisfies the following regret bound. See Appendix G.1 for a proof.

### v2

@@ 3.4 ExSUL: Exp3-IX for Online Selective Generation with Feedback Unlocking
Let $\ell_{t}(\cdot)\in[0,1]$ of the form (4). For any $T\in\mathbb{N}$ and finite hypotheses $\mathcal{{H}}$, Algorithm 1 provides the following regret bound with probability at least $1-\delta$, if $\eta=2\gamma=\sqrt{\frac{\ln|\mathcal{H}|}{T}}$:
$\textbf{Reg}_{T}\leq 4\sqrt{T\ln|\mathcal{H}|}+\left(1+\sqrt{\frac{T}{\ln|\mathcal{{H}}|}}\right)\ln\frac{2}{\delta}.$
Interestingly, despite of partial feedback, our algorithm achieves the same upper bound as EW with full feedback, i.e., $\sqrt{{T\ln|\mathcal{{H}}|}/{2}}$, up to a constant factor due to our loss estimator with feedback unlocking (7) for rich information, whereas Exp3-IX suffers from an additional factor of $\sqrt{|\mathcal{{H}}|}$, as in (10), due to its loss estimator with limited information.
@@ 4 Experiments
We empirically justify that ExSUL controls the FDR while maximizing selection efficiency under three diverse environments:(1) stochastic, (2) distribution-shift, and (3) interactive environments.
Datasets and Models. We use two datasets for stochastic and distribution-shift environments, $79\mathrm{K}$ Natural Question (NQ) (Kwiatkowski et al., 2019) and $93\mathrm{K}$ TriviaQA (Joshi et al., 2017) with two base models, GPT-3.5-turbo (OpenAI, 2023) and LLaMA3.1-8B-Instruct (Meta AI, 2024). We simulate distribution-shift environments by mixing two datasets in diverse ways and an interactive environment by generating interaction between two GPT-3.5-turbo models (see Section F.2 and F.3, respectively for details). To obtain feedback and self-consistency scores, we use GPT-3.5-turbo.
Scoring functions. We consider two scoring functions $f_{\texttt{std}}$ and $f_{\texttt{con}}$, introduced in Section B.1. In short, $f_{\texttt{std}}$ is a standard log-likelihood score defined as the conditional probability of an answer given a question. $f_{\texttt{con}}$ is a self-consistency score of an answer computed via entailment scores (from an entailment model) across samples. We use $f_{\texttt{con}}$ unless specified.
Methods. We consider two baselines, Exp3-IX-SG and No-SG, and one performance upper bound algorithm, EW-SG. In particular, (1) Exp3-IX-SG (Algorithm 6) is an online selective generator with partial feedback by adapting Exp3-IX (Neu, 2015) via our regret-to-FDR conversion lemma. (2) No-SG is a non-selective generator (i.e., $\tau_{t}=0$ for all $t$), to serve as a standard use of a generator without abstention. (3) EW-SG (Algorithm 5) is an online selective generator with full feedback. As it exploits full feedback, we use it as our empirical performance upper bound.
Metrics. We measure performance of ours along with comparing methods via the empirical FDR, i.e., $\textbf{FDR}_{t}$ and selection inefficiency, i.e., $\textbf{Ineff}_{t}$. Note that No-SG achieves zero selection inefficiency by definition, so we do not explicitly add in figures.
Parameters. We set $|\mathcal{{H}}|=1\mathrm{K}$ and $\lambda=T^{\nicefrac{{1}}{{4}}}$ by default. Also, a desired FDR level $\alpha$ and time horizon $T$ depend on tasks, where $\alpha$ is provided by users.
@@ 4.1 Stochastic Environment
Figure 3: Comparison of selective generation methods under a stochastic environment with LLaMA3.1-8B-Instruct as a generator on TriviaQA ($T=30\mathrm{K},\alpha=0.08$). The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples over 100100 random trials.
In Figure 3 and 8, we observe that ExSUL has clearly better convergence speed to the desired FDR compared to Exp3-IX-SG, since Exp3-IX-SG only observes $\ell_{t}(\tau_{t})$, requiring a much longer horizon $T$ (Figure 24). These results demonstrate that our partial feedback unlocking is beneficial in a faster convergence to the desired FDR, supporting the regret bound analysis discussed in Section 3.4.
In addition, Figure 22 shows that our algorithm effectively controls the FDR across different $\alpha$. Moreover, Figure 28 indicates that the inefficiency loss $a_{t}$ enables the algorithm to achieve efficient FDR guarantees. Note that algorithms with $f_{\texttt{con}}$ have better controllability over the FDR than those with $f_{\texttt{std}}$ (e.g., Figure 10 v.s. 12), suggesting that a well-calibrated scoring function can have a positive effect on convergence speed. See additional results in Section G.1.
4.2 Distribution-shift Environment
To evaluate the robustness of the methods, we consider three distribution-shift scenarios: (1) a single instantaneous change by concatenating two datasets, (2) frequent shifts by interleaving fixed-size chunks randomly drawn from each dataset, and (3) progressive shifts by sampling each example according to a linearly increasing mixing probability. See Appendix F.2 for details.
Interestingly, in Figure 4, 5, and 6, Exp3-IX-SG exhibits a sharp increase in FDR immediately following the distribution shift, highlighting its sensitivity to sudden changes in the data distribution, while the proposed ExSUL does not. In addition, Figure 23 shows that our algorithm still effectively controls the FDR across different $\alpha$ in the distribution shift environment. See additional results on varying generators and shift types in Section G.2 and a longer time horizon in Figure 25.
@@ 4.2 Distribution-shift Environment
Figure 4: Comparison of selective generation methods under a single distribution-shift environment with GPT-3.5-turbo as a generator ($T=30\mathrm{K},\alpha=0.1$), from TriviaQA to NQ. The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples with 100100 random trials.
@@
Figure 5: Comparison of selective generation methods under an alternating distribution-shift environment with GPT-3.5-turbo as a generator ($T=30\mathrm{K},\alpha=0.1$), alternating between TriviaQA and NQ, starting with TriviaQA. The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples with 100100 random trials.
@@
Figure 6: Comparison of selective generation methods under a gradual distribution-shift environment with GPT-3.5-turbo as a generator ($T=30\mathrm{K},\alpha=0.1$), from TriviaQA to NQ over time. The violin plots are drawn with randomly chosen $30\mathrm{K}$ samples with 100100 random trials.
4.3 Interactive Environment
We empirically demonstrate the efficacy of ExSUL in real-world interactive applications. In particular, we simulate an interactive environment by implementing a user-acting agent, a question-answering agent, and an evaluating agent by states, where these agents interact over multiple turns, and ExSUL decides whether to abstain from answering as shown in Figure 1.
In Figure 7, we observe that ExSUL consistently controls the FDR under $\alpha$, despite the instability of $\textbf{FDR}_{t}$ demonstrating the robustness of ExSUL under frequently shifting distributions over time. In Figure 1, ExSUL successfully abstains from incorrect answers, achieving the desired FDR level $\alpha$.
@@ 4.3 Interactive Environment
Figure 7: $\textbf{FDR}_{t}$ and $\textbf{Ineff}_{t}$ for ExSUL under an interactive environment with GPT-3.5-Turbo as a question-answering agent (T=12,500T=12{,}500, $\alpha=0.25$).
@@ 5 Conclusion
We propose a learning algorithm for online selective generation with partial feedback, which provides an FDR controllability guarantee, while allowing $\lambda$ to adjust the trade-off between the selection efficiency and convergence speed. To this end, we leverage the well-developed adversarial bandit problem. In particular, we reduce selective generation to adversarial bandits and extend the Exp3-IX algorithm for adversarial bandits by partial feedback unlocking, which addresses the lack of feedback information due to the nature of partial feedback. Finally, we introduce a novel Regret-to-FDR conversion lemma such that we can use any regret minimization algorithm, including our extended Exp3-IX for selective generation to control an FDR at a desired level. We theoretically and empirically justify our method by providing an efficient regret bound and evaluating diverse learning environments, including stochastic, distribution-shift, and interactive environments.
Limitations. The main limitation of our method is that our online selective generator selects a hypothesis independently of the observed input $\mathbf{x}_{t}$. As future work, extending to the contextual bandits, where the input is taken into account when pulling an arm, would be a promising direction.
@@ Acknowledgment
Appendix A LLM Usage
LLMs were used for minor grammar suggestions, and for assistance with code implementation.
Appendix B Preliminary
We introduce preliminaries on language generation, selective prediction, and adversarial bandits. To this end, let $\mathcal{{X}}$ be a set of inputs (e.g., examples or questions) and $\mathcal{{Y}}$ be a set of outputs (e.g., labels or answers).
B.1 Language Generation
[…]
B.2 Selective Prediction
Selective prediction (Geifman & El-Yaniv, 2017) and generation (Lee et al., 2024) provide certified control over the risk of incorrect predictions by abstaining answers (saying IDK) when the prediction is uncertain. Given a predictor $\hat{\mathbf{y}}:\mathcal{{X}}\to\mathcal{{Y}}$, a selective predictor $\hat{{S}}:\mathcal{{X}}\to\mathcal{{Y}}\cup\{\texttt{IDK}\}$ abstains from returning $\hat{\mathbf{y}}(\mathbf{x})$ if a selection function $\hat{{s}}:\mathcal{{X}}\times\mathcal{{Y}}\to\{0,1\}$ deems the prediction uncertain, i.e., $\hat{{S}}(\mathbf{x})\coloneqq\begin{cases}\hat{\mathbf{y}}(\mathbf{x})&\text{if~}\hat{{s}}(\mathbf{x},\hat{\mathbf{y}}(\mathbf{x}))=1\\ \texttt{IDK}&\text{otherwise}\end{cases}.$ In learning, the selection function $\hat{{s}}$ is chosen to possibly satisfy a desired level of a false discovery rate i.e., $\mathbbm{P}(\hat{{S}}(\mathbf{x})\neq_{E}\mathbf{y}\mid\hat{{S}}(\mathbf{x})\neq\texttt{IDK})$ over the i.i.d. samples of $(\mathbf{x},\mathbf{y})$. In selective classification (Geifman & El-Yaniv, 2017), the equality relation $\neq_{E}$ is usually the standard exact matching, which is extended to capture semantic correctness via entailment relation in selective generation (Lee et al., 2024). In contrast to the standard stochastic setup as in the previous literature, we consider online learning under partial feedback.
B.3 Regret Minimization
Sequential prediction in multi-armed bandit problems is formulated as an interactive game between a learner and an adversary (also called an environment). In particular, for each time $t\in\{1,\dots,T\}$, a learner selects an arm $\tau_{t}\in\mathcal{{H}}$ among a set of arms $\mathcal{{H}}$ and an adversary selects a loss function $\ell_{t}:\mathcal{{H}}\to\mathbb{R}_{\geq 0}$, where a loss associated to an arm $\tau_{t}$ is denoted by $\ell_{t}(\tau_{t})$ and provided to the learner as partial feedback (i.e., only the loss for the selected arm is provided). Note that this interactive game often describes in rewards, but we consider losses, instead of rewards.
The main goal of this interactive game is to find a learner whose performance is as good as the best learner in hindsight. Here, the performance is measured by regret, following standard bandits literatures (Bubeck et al., 2012), i.e.,
$\textbf{Reg}_{T}\coloneqq\sum_{t=1}^{T}\ell_{t}(\tau_{t})-\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\ell_{t}(\tau).$ (8)
Thus, the the goal becomes to find a learner that minimizes the regret such that an achieved regret bound is sublinear in $T$. Here, we say that the adversary is oblivious if the sequence of loss functions are chosen regardless of learner’s previous selections (or chosen in advance before the game), and the adversary is adaptive if the loss function at time $t$, i.e., $\ell_{t}$, is drawn a distribution that depends on the previously chosen arms, i.e., $\tau_{1},\dots,\tau_{t-1}$. In this paper, we consider the adaptive adversary which is more stronger than the oblivious one. Moreover, under the adaptive adversary setup, a deterministic learner cannot win the game (i.e., it does not achieve a sub-linear regret bound), so we consider a randomized learner. This implies that $\tau_{t}$ and $\ell_{t}$ are two sources of randomness, so $\textbf{Reg}_{T}$ is a random variable. Thus, we consider the high-probable regret bound or a bound of an expected regret $\mathbbm{E}\textbf{Reg}_{T}$. Note that we only consider the learner’s randomization in the analysis, since our bounds are derived to hold for any adversary (Bubeck et al., 2012; Neu, 2015).
In the following section, we review one regret minimization algorithm for bandits with an adaptive adversary, called Exp3-IX, and see Appendix E for other algorithms.
B.4 Adversarial Bandits and Exp3-IX Algorithm
Traditional online learning typically assumes full feedback, i.e., loss for any hypothesis is computable. However, practical applications may only have partial feedback, i.e., loss for a chosen hypothesis is only given. We model this via the adversarial multi-armed bandit problem. The regret is as in online learning, but the learner only receives for a chosen arm, while the sequence $\ell_{t}(\tau_{t})$ is chosen by an oblivious or adaptive adversary that may condition on the history of chosen arms but not on the learner’s fresh randomness at time $t$.
To obtain the high-probability control of the $\textbf{Reg}_{T}$ over an adaptive adversary, Exponential-weights for Exploration and Exploitation with Implicit Exploration (Exp3-IX)(Neu, 2015) is a natural choice. It employs implicit-exploration (IX) loss estimates within an exponential-weights scheme and does not require explicit mixing with the uniform distribution like Exp3.P (Auer et al., 2002). In particular, Exp3-IX uses the variance-reduced loss estimator for each arm $\tau\in\mathcal{{H}}$ constructed from the observed loss of the chosen arm $\tau_{t}$, i.e.,
$\tilde{\ell}_{t}(\tau\mid\{\tau_{t}\})\coloneqq\frac{\ell_{t}(\tau)}{\gamma_{t}+p_{t}(\tau)}\cdot\mathbbm{1}(\tau_{t}=\tau)\in[0,\infty).$ (9)
where $\gamma_{t}>0$ is properly selected by an analysis; thus by using (9) in $\textbf{Reg}_{T}$, with non-increasing learning rate $\eta_{t}$, we have the following regret with high probability for Exp3-IX . See AppendixH.2 for a complete proof.